Comment epi2 SNF 2010-04-15 15:11:07: Update bulb failure

eenriquez at snf.stanford.edu eenriquez at snf.stanford.edu
Thu Apr 15 15:11:07 PDT 2010


Unable to find the "bad" bulb.  Ohmmeter readings for all the lamp loops were good (1.8 ohm).  Started checking the lamp sockets. No obviously bad socket,  need to check them one at a time.




More information about the epi2-pcs mailing list